# Moduli of bundles and semiorthogonal decomposition

### v1

@@ Abstract.
In this paper we construct semiorthogonal decompositions of moduli of principal bundles on a curve into its symmetric powers, for both the moduli stack of all $G$-bundles and the coarse moduli space of semistable $G$-bundles. The essential ingredients in the proof include Borel-Weil-Bott theory for loop groups, highest weight structure of current group representation, derived $\Theta$-stratification and local-global compatibility of Kac-Moody localization.
@@ 1. Introduction
Moduli theory provides a rich source of higher-dimensional Fano varieties: let $X$ be a smooth projective curve of genus $g>1$, $L$ be a line bundle on $X$, then the coarse moduli space [^1] We will always use Roman type for coarse moduli space and Italic type for moduli stack. $Bun_{n}^{L,ss}$ of semistable vector bundles on $X$ with rank $n$ and determinant $L$, are Fano varieties of dimension $(g-1)(n^{2}-1)$ (smooth if $\delta=(n,\deg L)=1$ and with Gorenstein canonical singularity in general [beauville1998picard]). Analogous construction exists for any semisimple algebraic groups $G$. In low rank examples, they are known to admit a motivic decomposition into products of symmetric powers $\mathrm{Sym}^{\lambda_{1}}X\times\cdots\times\mathrm{Sym}^{\lambda_{n-1}}X$, and they are believed to lift to a semiorthogonal decomposition. [lee2018remarks, lee2021symmetric]
@@
The derived categories and semiorthogonal decompositions are well-behaved (i.e. saturated) in the smooth case, and in the singular case, we can replace $\mathrm{D}^{b}(Bun_{n}^{L,ss})$ by a categorical/non-commutative resolution of singularity $\mathrm{D}^{b}(Bun_{n}^{L,ss})^{\sim}$. There are natural constructions from the rigidified moduli stack $\mathrm{D}^{b}(\mathrm{Bun}_{n}^{L,ss}/\mathrm{B}\mu_{n})$ (and similarly for $\mathrm{D}^{b}(Bun_{n}^{d,ss})$).
@@
where $\theta$ is the theta line bundle, integers $(k,\lambda_{1},\cdots,\lambda_{r})$ satisfy $0\leq k\leq\delta$, $(g-1)\rho-\lambda$ lies in the cone generated by positive roots, and lies in the interior if $k=\delta$. Here $\rho$ is half the sum of positive roots, $\lambda=\sum_{i}\lambda_{i}\omega_{i}$ is the dominant weight with coefficients $\lambda_{i}$, $\delta=\frac{h^{\vee}}{\mathrm{ord}(\xi)}$, $h^{\vee}$ is the dual Coxeter number, and $o$ is the order of the central character of the determinant line bundle. $\cdots$ is the semiorthogonal complement of the previous subcategories, and $\mathrm{D}^{b}({Bun}_{r}^{L,ss})^{\sim}$ is a categorical resolution of singularity of ${Bun}_{G}^{\xi,ss}$ (which is $\mathrm{D}^{b}({Bun}_{r}^{L,ss})$ for smooth ${Bun}_{G}^{\xi,ss}$)
@@
This celebrated conjecture, raised independently by Belmans-Galkin-Mukhopadhyay and Narasimhan, was proved [^2] None of the two papers prove $\cdots=0$ half of the conjecture, it was proved later by Tevelev in [tevelev2023braid]. independently by Tevelev-Torres[tevelev2021narasimhan] and Xu-Yau[xu2021semiorthogonal]. See also [lee2023derived] for following works. Our previous method does not seem to yield a complete decomposition for $r>3$, although it provides partial decompositions as discussed above.
@@
$\mathrm{D}^{b}({\mathrm{Bun}}_{n}^{L,ss})=\langle\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda_{1}}X\times\cdots\times\mathrm{Sym}^{\lambda_{n-1}}X)\rangle$
where $\mathcal{L}$ is the fundamental determinant line bundle, $k\in\mathbb{Z}$ and $\lambda\in\mathbb{N}^{r}$.
The open substack $\mathrm{Bun}_{G}^{\xi,ss}\longrightarrow\mathrm{Bun}_{G}^{\xi}$ is the semistable locus determined by the Harder-Narasimhan filtration, and the machinery of derived $\Theta$-stratification developed by Halpern-Leistner allows us to relate their derived categories. $\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi,ss})\longrightarrow\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})$ is a full subcategory which satisfies certain restrictions of weights on the polystable locus. (This is how the finite subsets arise.) So Theorem 3 is reduced to the following:
@@
Here we are taking the restricted product, the reader is referred to Section 2 for details. We prove that for nontrivial $\xi$ we need to twist the right-hand side by the outer-automorphism of $G(K_{X})$ corresponding to $\xi$:
@@
It is enlightening to notice that $\mathrm{B}G(\mathcal{O})$ is the moduli of $G$ bundle on $D=\mathrm{Spec}(\mathcal{O})$ while $\mathrm{Gr}_{G}$ is the moduli of $G$ bundle on $D=\mathrm{Spec}(\mathcal{O})$ equipped with a trivialization on $D^{\circ}=\mathrm{Spec}(\mathcal{K})$, thus Theorems 6-7 can be viewed as local analogues of Theorem 3 where we take $X=D$ with only ceveat that in the non compactly-supported setting the determinant line bundle it not defined. Theorem 6 follows from the highest weight structure of $\mathrm{Rep}(G(\mathcal{O}))$ and Theorem 7 follows from the affine Borel-Weil-Bott theorem, the triviality of loop group representation at level 0 and Theorem 6.
@@
This method was applied in an earlier version of the paper [xu2023moduli] to prove the main theorems for type A (and straightforwardly generalizes to other classical groups). The machinery of global Weyl modules [khoroshkin2013highest, bennett2014macdonald, chari2015bgg] furnishes a parallel proof that avoids explicit use of Schur–Weyl duality and the stacky symmetric-power decomposition. These modules are the standard objects in the highest weight category $\mathrm{Rep}(G(\mathcal{O}))$, whose highest weight structures are induced from $\mathrm{Rep}\,G$. It is known that this construction is functorial, and we show that it can be glued from affine charts to a globally defined functor on products of symmetric powers. This is precisely the functor needed for the semiorthogonal decomposition; once it is in place, semiorthogonality follows from Borel–Weil–Bott for loop groups together with local calculations for global Weyl modules.
@@ 1.1. Acknowledgements
The author is grateful to S.-T. Yau for guidance and support over the years; to L. Chen and Y. Fu for teaching him geometric representation theory; and to D. Auroux, R. Fujita, D. Halpern-Leistner, H.-B. Moon, and D. Ben-Zvi for helpful discussions.
@@ 2. Moduli of twisted bundles
Throughout this paper, we fix an algebraically closed base field of characteristic 0.333All of our constructions are canonical and Galois-equivariant, hence straightforwardly descend to non-closed fields. Let $G$ be a simple algebraic group, $X$ be a smooth projective curve and $\mathrm{Bun}_{G}=\mathrm{Map}(X,\mathrm{B}G)$ be the moduli stack of $G$-bundles over $X$. Let $Z=Z$ be the center of $G$, $X,Y$ be the weight and root lattices, the $Z$ gerbes are classifield by $\mathrm{Map}(X,\mathrm{B}^{2}Z)$, by Poincare duality this is isomorphic to $\mathrm{Hom}(\mathrm{Map}(X,\mathrm{Hom}(Z,\mathbb{G}_{m})),\mathbb{Q}/\mathbb{Z})$. Note that $\mathrm{Hom}(Z,\mathbb{G}_{m})\simeq Q$ is the quotient of root lattice of $G$ by weight lattice of $G$, we have a natural $\mathbb{Q}$-valued inner product on the weight lattice, such that root lattice is dual to weight lattice, so we get a $\mathbb{Q}/\mathbb{Z}$ valued bilinear pairing on their quotient $Q$ and this is a perfect pairing. Hence it induces a natural isomorphism $Q\simeq\mathrm{Hom}(Q,\mathbb{Q}/\mathbb{Z})$, and the $Z$ gerbes are classified by the homology of $X$ with coefficient $Q$, i.e. the isomorphism classes are classified by $Q$, the automorphisms are $H_{1}(X,Q)$ and the $2$-automorphism are $H_{2}(X,Q)$. Let $G^{ad}=G/Z$ be the adjoint group of $G$, given a $Z$-gerbe $\xi$, we define the moduli of $\xi$-twisted $G$ bundle $\mathrm{Bun}^{\xi}_{G}$ to be the fiber of $\mathrm{Map}(X,\mathrm{B}G^{ad})\longrightarrow\mathrm{Map}(X,\mathrm{B}^{2}Z)$ at $\xi$, which recover $\mathrm{Bun}_{G}$ when $\xi$ is the trivial gerbe. [^4] Note that $\mathrm{Map}(X,\mathrm{B}^{2}Z)$ is a finite abelian group stack with nontrivial $\pi_{0},\pi_{1},\pi_{2}$ and we cannot replace it by $H^{2}(X,Z)$. Let $\mathrm{Bun}_{G}^{\xi,s}$ (resp, $\mathrm{Bun}_{G}^{\xi,ss}$) be the open substack of stable (resp, semistable) twisted bundles. For $G=\mathrm{SL}_{n}$, we have an equivalent description $\mathrm{Bun}^{\xi}_{G}\simeq\mathrm{Bun}_{n}^{L}$ as the moduli stack of rank $n$ vector bundles $E$ with an isomorphism $\det E\simeq L$ for some fixed $L$ whose degree equals to $\xi$ in $Q\simeq\mathbb{Z}/(r+1)\mathbb{Z}$. [^5] Throughout the paper we use the étale topology on $X$.. The equivalence follows from this Cartesian diagram of classifying stacks:
@@ 5. Reduction to coarse moduli
The basic idea is to apply the excision sequence to the inner homomorphism $H=\mathcal{H}om(E,F)\in\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})$:
@@
Let us study when there first term $\Gamma_{\mathrm{Bun}_{G}^{\xi,us}}(\mathrm{Bun}_{G}^{\xi},H)=0$. We need to analyze each unstable stratum individually. Note that the Harder-Narasimhan stratification naturally gives rise to maps $\mathrm{Bun}_{G}^{\xi,\delta,ps}\overset{\sigma_{\delta}}{\underset{\pi_{\delta}}{\rightleftarrows}}\mathrm{Bun}_{G}^{\xi,\delta,us}\overset{i_{\delta}}{\rightarrow}\mathrm{Bun}_{G}^{\xi}$ for each Harder-Narasimhan stratum labelled by the parabolic weight $\delta$. On the polystable loci $\mathrm{Bun}_{G}^{\xi,\delta,ps}$ there is a natural $\mathrm{B}\mathbb{G}_{m}$ action coming from $\xi$ with respect to which the weights of $\pi_{\delta*}(\mathcal{O}_{\mathrm{Bun}_{G}^{\xi,\delta,us}})$ are non-negative, so a sufficient condition for $\Gamma_{\mathrm{Bun}_{G}^{\xi,us}}(\mathrm{Bun}_{G}^{\xi},H)=0$ would be that the $\mathbb{G}_{m}$ weights of $\sigma_{\delta}^{*}i_{\delta}^{!}H$ to be strictly positive. It’s straightforward calculation that for $F\in\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda}X)$ and $E\in\mathcal{L}^{\otimes l}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\mu}X)$, the lowest $\mathbb{G}_{m}$ weight is at least $2(g-1)\rho-(\lambda^{*}+\mu),\delta>+(h^{\vee}-(k-l))|\delta|^{2}$, where * means the Dynkin involution on the weight lattice coming from the duality in $\mathcal{H}om(E,F)=E^{*}\otimes F$ , and $h^{\vee}=h^{\vee}_{G}$ is the dual Coxeter number of $G$. So we see that so long as $2(g-1)\rho-(\lambda^{*}+\mu)$ lies in the positive cone (i.e. cone spanned by the positive roots) and $k-l\leq h^{\vee}$, and they do not lie on the boundary of either inequalities, the lowest weight is strictly positive, and we have $\Gamma_{\mathrm{Bun}_{G}^{\xi,us}}(\mathrm{Bun}_{G}^{\xi},H)=0$. From this we directly see that the composition $\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda}X)\longrightarrow\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})\longrightarrow\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi,ss})$ is fully faithful assuming $2(g-1)\rho-(\lambda^{*}+\lambda)$ lies in the positive cone, and moreover the semiorthogonal relation of $\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda}X)$ and $\mathcal{L}^{\otimes l}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\mu}X)$ in $\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})$ is preserved in $\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi,ss})$ if $2(g-1)\rho-(\lambda^{*}+\mu)$ lies in the positive cone. Hence we get Theorem 13. Note that the components of this decomposition are weight subcategories of the $\mathrm{B}Z$ actions, we may directly take the ones with trivial $\mathrm{B}Z$ action, which are exactly those which satisfy the central character of $\mathcal{L}^{\otimes k}$ cancels with the central character of the representation with highest weight $\lambda$.

### v2

@@ Abstract.
In this paper we construct semiorthogonal decompositions of moduli of principal bundles on a curve into its symmetric powers, for both the moduli stack of all $G$-bundles and the coarse moduli space of semistable $G$-bundles. The essential ingredients in the proof include Borel-Weil-Bott theory for loop groups, highest weight structure of current group representation, derived $\Theta$-stratification and local-global compatibility of Kac-Moody localization. For the full moduli stack of all $G$-bundles the decomposition we construct is complete, while for the coarse moduli space of semistable $G$-bundles it is not clear in full generality yet.
@@ 1. Introduction
Moduli theory provides a rich source of higher-dimensional Fano varieties: let $X$ be a smooth projective curve of genus $g>1$, $L$ be a line bundle on $X$, then the coarse moduli space [^1] We will always use Roman type for coarse moduli space and Italic type for moduli stack. $Bun_{n}^{L,ss}$ of semistable vector bundles on $X$ with rank $n$ and determinant $L$, are Fano varieties of dimension $(g-1)(n^{2}-1)$ (smooth if $\delta=(n,\deg L)=1$ and with Gorenstein canonical singularity in general [beauville1998picard]). Analogous construction exists for any semisimple algebraic groups $G$. In low rank examples, they are known to admit a motivic decomposition into products of symmetric powers $\mathrm{Sym}^{\lambda_{1}}X\times\cdots\times\mathrm{Sym}^{\lambda_{n-1}}X$. They are believed to lift to a semiorthogonal decomposition [lee2018remarks, lee2021symmetric], although we only have partial decompositions yet..
@@
With our choice of Fourier-Mukai functor defined by the global Weyl modules, the order of blocks is that less dominant $\lambda$ maps to more dominant $\lambda$, and smaller $k$ maps to larger $k$. The derived categories and semiorthogonal decompositions are well-behaved (i.e. saturated) in the smooth case, and in the singular case, we can replace $\mathrm{D}^{b}(Bun_{n}^{L,ss})$ by a categorical/non-commutative resolution of singularity $\mathrm{D}^{b}(Bun_{n}^{L,ss})^{\sim}$  [^2] It would be natural to compare with the quasi-BPS categories [davison2023bps, puadurariu2025quasi]. For $G=SL_{2}$, the quasi-BPS category was completely understood in [sink2025noncommutative]. We thank J. Tevelev for this comment.. There are natural constructions from the rigidified moduli stack $\mathrm{D}^{b}(\mathrm{Bun}_{n}^{L,ss}/\mathrm{B}\mu_{n})$ (and similarly for $\mathrm{D}^{b}(Bun_{n}^{d,ss})$).
@@
where $\theta$ is the theta line bundle, integers $(k,\lambda_{1},\cdots,\lambda_{r})$ satisfy $0\leq k<2\delta$, $(g-1)\rho-\lambda$ lies in the cone generated by positive roots, and lies in the interior if $k=\delta$. Here $\rho$ is half the sum of positive roots, $\lambda=\sum_{i}\lambda_{i}\omega_{i}$ is the dominant weight with coefficients $\lambda_{i}$, $\delta=\frac{h^{\vee}}{\mathrm{ord}(\xi)}$, $h^{\vee}$ is the dual Coxeter number. $\cdots$ is the semiorthogonal complement of the previous subcategories, and $\mathrm{D}^{b}({Bun}_{r}^{L,ss})^{\sim}$ is a categorical resolution of singularity of ${Bun}_{G}^{\xi,ss}$ (which is $\mathrm{D}^{b}({Bun}_{r}^{L,ss})$ for smooth ${Bun}_{G}^{\xi,ss}$)
@@
This celebrated conjecture, raised independently by Belmans-Galkin-Mukhopadhyay and Narasimhan, was proved [^3] None of the two papers prove $\cdots=0$ half of the conjecture, it was proved later by Tevelev in [tevelev2023braid]. independently by Tevelev-Torres[tevelev2024bgmn] and Xu-Yau[xu2021semiorthogonal]. See also [lee2023derived] for following works. Our previous method does not seem to yield a complete decomposition for $r>3$, although it provides partial decompositions as discussed above.
@@
$\mathrm{D}^{b}({\mathrm{Bun}}_{G}^{\xi,ss})=\langle\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda_{1}}X\times\cdots\times\mathrm{Sym}^{\lambda_{n-1}}X),\cdots\rangle$
where $\mathcal{L}$ is the fundamental determinant line bundle, $k\in\mathbb{Z}$ and $\lambda\in\mathbb{N}^{r}$ satisfy $0\leq k<2h^{\vee}$, $\sum_{i\leq l}(\lambda_{i}-(g-1))i(n-l)+\sum_{i>l}(\lambda_{i}-(g-1))l(n-i)\leq 0$ for all $0<l<n$, and when $k\geq h^{\vee}$ all other upper bounds are strict. $\cdots$ is the semiorthogonal complement of the previous subcategories.
The open substack $\mathrm{Bun}_{G}^{\xi,ss}\longrightarrow\mathrm{Bun}_{G}^{\xi}$ is the semistable locus determined by the Harder-Narasimhan filtration, and the machinery of derived $\Theta$-stratification developed by Halpern-Leistner allows us to relate their derived categories. $\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi,ss})\longrightarrow\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})$ is a full subcategory which satisfies certain restrictions of weights on the unstable loci. (This is how the finite subsets arise.) So Theorem 3 is reduced to the following:
@@
Here $\mathcal{O}_{x}\simeq k[[t]]$, $\mathcal{K}_{x}\simeq k((t))$, and $K_{X}$ is the function field of $X$. We are taking the restricted product, the reader is referred to Section 2 for details. We prove that for nontrivial $\xi$ we need to twist the right-hand side by the outer-automorphism of $G(K_{X})$ corresponding to $\xi$:
@@
It is enlightening to notice that $\mathrm{B}G(\mathcal{O})$ is the moduli of $G$ bundle on $D=\mathrm{Spec}(\mathcal{O})$ while $\mathrm{Gr}_{G}$ is the moduli of $G$ bundle on $D=\mathrm{Spec}(\mathcal{O})$ equipped with a trivialization on $D^{\circ}=\mathrm{Spec}(\mathcal{K})$, thus Theorems 7,8 can be viewed as local analogues of Theorem 5 where we take $X=D$ with only ceveat that in the non compactly-supported setting the determinant line bundle it not defined. Theorem 7 follows from the highest weight structure of $\mathrm{Rep}(G(\mathcal{O}))$ and Theorem 8 follows from the affine Borel-Weil-Bott theorem, the triviality of loop group representation at level 0 and Theorem 6.
@@
This method was applied in an earlier version of the current paper [xu2023moduli] to prove the main theorems for type A (and straightforwardly generalizes to other classical groups). The machinery of global Weyl modules [khoroshkin2013highest, bennett2014macdonald, chari2015bgg] furnishes a parallel proof that avoids explicit use of Schur–Weyl duality and the stacky symmetric-power decomposition. These modules are the standard objects in the highest weight category $\mathrm{Rep}(G(\mathcal{O}))$, whose highest weight structures are induced from $\mathrm{Rep}\,G$. It is known that this construction is functorial, and we show that it can be glued from affine charts to a globally defined functor on products of symmetric powers. This is precisely the functor needed for the semiorthogonal decomposition; once it is in place, semiorthogonality follows from Borel–Weil–Bott for loop groups together with local calculations for global Weyl modules.
@@ 1.1. Acknowledgements
The author is grateful to S.-T. Yau for guidance and support over the years; to L. Chen and Y. Fu for teaching him geometric representation theory; to J. Tevelev for many helpful comments on the first draft, and to D. Auroux, R. Fujita, D. Halpern-Leistner, H.-B. Moon, and D. Ben-Zvi for helpful discussions.
@@ 2. Moduli of twisted bundles
Throughout this paper, we fix an algebraically closed base field of characteristic 0.444All of our constructions are canonical and Galois-equivariant, hence straightforwardly descend to non-closed fields. Let $G$ be a simple algebraic group, $X$ be a smooth projective curve and $\mathrm{Bun}_{G}=\mathrm{Map}(X,\mathrm{B}G)$ be the moduli stack of $G$-bundles over $X$. Let $Z=Z$ be the center of $G$, $X,Y$ be the weight and root lattices, the $Z$ gerbes are classifield by $\mathrm{Map}(X,\mathrm{B}^{2}Z)$, by Poincare duality this is isomorphic to $\mathrm{Hom}(\mathrm{Map}(X,\mathrm{Hom}(Z,\mathbb{G}_{m})),\mathbb{Q}/\mathbb{Z})$. Note that $\mathrm{Hom}(Z,\mathbb{G}_{m})\simeq Q$ is the quotient of root lattice of $G$ by weight lattice of $G$, we have a natural $\mathbb{Q}$-valued inner product on the weight lattice, such that root lattice is dual to weight lattice, so we get a $\mathbb{Q}/\mathbb{Z}$ valued bilinear pairing on their quotient $Q$ and this is a perfect pairing. Hence it induces a natural isomorphism $Q\simeq\mathrm{Hom}(Q,\mathbb{Q}/\mathbb{Z})$, and the $Z$ gerbes are classified by the homology of $X$ with coefficient $Q$, i.e. the isomorphism classes are classified by $Q$, the automorphisms are $H_{1}(X,Q)$ and the $2$-automorphism are $H_{2}(X,Q)$. Let $G^{ad}=G/Z$ be the adjoint group of $G$, given a $Z$-gerbe $\xi$, we define the moduli of $\xi$-twisted $G$ bundle $\mathrm{Bun}^{\xi}_{G}$ to be the fiber of $\mathrm{Map}(X,\mathrm{B}G^{ad})\longrightarrow\mathrm{Map}(X,\mathrm{B}^{2}Z)$ at $\xi$, which recover $\mathrm{Bun}_{G}$ when $\xi$ is the trivial gerbe. [^5] Note that $\mathrm{Map}(X,\mathrm{B}^{2}Z)$ is a finite abelian group stack with nontrivial $\pi_{0},\pi_{1},\pi_{2}$ and we cannot replace it by $H^{2}(X,Z)$. Let $\mathrm{Bun}_{G}^{\xi,s}$ (resp, $\mathrm{Bun}_{G}^{\xi,ss}$) be the open substack of stable (resp, semistable) twisted bundles. For $G=\mathrm{SL}_{n}$, we have an equivalent description $\mathrm{Bun}^{\xi}_{G}\simeq\mathrm{Bun}_{n}^{L}$ as the moduli stack of rank $n$ vector bundles $E$ with an isomorphism $\det E\simeq L$ for some fixed $L$ whose degree equals to $\xi$ in $Q\simeq\mathbb{Z}/(r+1)\mathbb{Z}$. [^6] Throughout the paper we use the étale topology on $X$.. The equivalence follows from this Cartesian diagram of classifying stacks:
@@ 5. Reduction to coarse moduli
Following [2020arXiv201001127H], the basic idea is to apply the excision sequence to the internal homomorphism $H=\mathcal{H}om(E,F)\in\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})$:
@@
Let us study when there first term $\Gamma_{\mathrm{Bun}_{G}^{\xi,us}}(\mathrm{Bun}_{G}^{\xi},H)=0$. We need to analyze each unstable stratum individually. Note that the Harder-Narasimhan stratification naturally gives rise to maps $Z_{G}^{\xi,\delta}\overset{\sigma_{\delta}}{\underset{\pi_{\delta}}{\rightleftarrows}}\mathrm{Bun}_{G}^{\xi,\delta,us}\overset{i_{\delta}}{\rightarrow}\mathrm{Bun}_{G}^{\xi}$ for each Harder-Narasimhan stratum labelled by the parabolic weight $\delta$. On its center $Z_{G}^{\xi,\delta}$ there is a natural $\mathrm{B}\mathbb{G}_{m}$ action [^15] GG action on a category gives $G$ action on the collection of objects, and $\mathrm{B}G$ action on a category gives $G$ action on each space of morphisms. coming from $\xi$ with respect to which the weights of $\pi_{\delta*}(\mathcal{O}_{\mathrm{Bun}_{G}^{\xi,\delta,us}})$ are non-negative, so a sufficient condition for $\Gamma_{\mathrm{Bun}_{G}^{\xi,us}}(\mathrm{Bun}_{G}^{\xi},H)=0$ would be that the $\mathbb{G}_{m}$ weights of $\sigma_{\delta}^{*}i_{\delta}^{!}H$ to be strictly positive. It’s straightforward calculation that for $F\in\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda}X)$ and $E\in\mathcal{L}^{\otimes l}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\mu}X)$, the lowest $\mathbb{G}_{m}$ weight is at least $2(g-1)\rho-(\lambda^{*}+\mu),\delta>+(h^{\vee}-(k-l))|\delta|^{2}$, where * means the Dynkin involution on the weight lattice coming from the duality in $\mathcal{H}om(E,F)=E^{*}\otimes F$ , and $h^{\vee}=h^{\vee}_{G}$ is the dual Coxeter number of $G$. So we see that so long as $2(g-1)\rho-(\lambda^{*}+\mu)$ lies in the positive cone (i.e. cone spanned by the positive roots) and $k-l\leq h^{\vee}$, and they do not lie on the boundary of either inequalities, the lowest weight is strictly positive, and we have $\Gamma_{\mathrm{Bun}_{G}^{\xi,us}}(\mathrm{Bun}_{G}^{\xi},H)=0$. From this we directly see that the composition $\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda}X)\longrightarrow\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})\longrightarrow\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi,ss})$ is fully faithful assuming $2(g-1)\rho-(\lambda^{*}+\lambda)$ lies in the positive cone, and moreover the semiorthogonal relation of $\mathcal{L}^{\otimes k}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\lambda}X)$ and $\mathcal{L}^{\otimes l}\otimes\mathrm{D}^{b}(\mathrm{Sym}^{\mu}X)$ in $\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi})$ is preserved in $\mathrm{D}^{b}(\mathrm{Bun}_{G}^{\xi,ss})$ if $2(g-1)\rho-(\lambda^{*}+\mu)$ lies in the positive cone. Hence we get Theorem 13. Note that the components of this decomposition are weight subcategories of the $\mathrm{B}Z$ actions, we may directly take the ones with trivial $\mathrm{B}Z$ action, which are exactly those which satisfy the central character of $\mathcal{L}^{\otimes k}$ cancels with the central character of the representation with highest weight $\lambda$.